OpenStudy (anonymous):

Use the order of operations to simplify the expression. twenty minus sixteen plus 8 times 2

OpenStudy (anonymous):

20-16+8(2)=20

OpenStudy (anonymous):

20-16+16=20

OpenStudy (anonymous):

4+16=20

OpenStudy (anonymous):

20=20